# Parameterized Controls and Isomorphic Representations of G-Expectations

### v1

@@ 5.2 Asymmetric Martingales: Favorable Gambles with Worst-Case Fairness
    For all $\theta\in\mathcal{A}_{s,t}^{\Theta}$, $M_{t}^{\theta}$ is a supermartingale under $\mathbb{P}$:
    $\mathbb{E}_{\mathbb{P}}[M_{t}^{\theta}|\mathcal{F}_{s}]\leq M_{s}^{\theta}\quad\forall s\leq t$
@@
    There exists a worst-case scenario $\theta^{*}\in\Theta$ such that:
    $\mathbb{E}_{\mathbb{P}}[M_{t}^{\theta^{*}}|\mathcal{F}_{s}]=M_{s}^{\theta^{*}}$
5.3 Key Properties
[…]
@@ 6 Conclusion
This paper has established a fundamental and rigorous connection between the abstract theory of $G$-expectation and the concrete framework of parameterized stochastic control. By introducing the concept of controlled processes $\{B_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$, we have constructed a complete isomorphism $\Phi$ that maps the nonlinear $G$-expectation space onto a family of classical probability spaces indexed by volatility controls.
The core contribution of this work is threefold:
  1. 1.
    Unified Representation: We demonstrated that $G$-Brownian motion and $G$-SDE solutions admit a natural representation as suprema over parameterized classical processes, providing an intuitive and computationally tractable interpretation of model uncertainty.
  2. 2.
    Algebraic and Analytic Isomorphism: The mapping $\Phi$ was shown to preserve all essential operations – addition, multiplication, function composition, and limit transitions – thereby maintaining the complete mathematical structure while transitioning between nonlinear and classical frameworks.
  3. 3.
    Martingale Classification: Our framework provided a clear distinction between symmetric $G$-martingales (fair in all volatility scenarios) and asymmetric $G$-martingales (favorable but fair only in worst-case scenarios), revealing the fundamental structure of stochastic processes under volatility uncertainty.
The parameterized control perspective developed herein offers more than theoretical elegance; it provides a practical computational methodology for working with $G$-expectations by reducing complex nonlinear problems to families of more manageable classical problems. This approach maintains full mathematical rigor while offering significant advantages for both theoretical analysis and numerical implementation.
Looking forward, this isomorphism establishes a solid foundation for further research in nonlinear stochastic calculus, with particular promise for applications in mathematical finance where volatility uncertainty plays a crucial role in pricing and risk management. The framework naturally extends to more complex stochastic processes and offers a new paradigm for understanding model uncertainty in quantitative economics.

### v7

@@ 8.1 Unification via Linear Decomposition
    Linear Execution Phase: Stochastic analysis operations—such as proving inequalities, solving SDEs, and establishing estimates—are performed within the familiar, linear environment of classical probability theory for each fixed control parameter $\theta\in\mathcal{A}^{\Theta}$.
@@
    Uniform Synthesis Phase: The results from the parameterized family are synthesized into the robust $G$-framework through the single operation of taking the supremum over the parameter space $\Theta$, provided uniform estimates are established.
8.1.1 The Unification of Fundamental Inequalities
We demonstrate the utility of this decomposition principle by unifying and simplifying the proofs of fundamental inequalities within the $G$-framework. The general procedure consists of three steps: Translation, Linear Execution, and Uniform Synthesis.
Hölder Inequality Parametric Derivation:
  1. 1.
    Translation: For $X,Y\in L_{G}^{p}$, by definition of the homomorphism: $\hat{\mathbb{E}}[|XY|]=\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}Y^{\theta}\right|\right]$.
  2. 2.
    Linear Execution: For each fixed $\theta$, apply the classical Hölder inequality: $\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}Y^{\theta}\right|\right]\leq\left(\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}\right|^{p}\right]\right)^{1/p}\left(\mathbb{E}_{\mathbb{P}}\left[\left|Y^{\theta}\right|^{q}\right]\right)^{1/q}$, where $1/p+1/q=1$.
  3. 3.
    Uniform Synthesis: Taking the supremum over $\theta$ on both sides. Using the property $\sup_{\theta}(f(\theta)g(\theta))\leq\sup_{\theta}f(\theta)\sup_{\theta}g(\theta)$ for non-negative functions, we obtain the $G$-Hölder inequality:
    $\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}Y^{\theta}\right|\right]\leq\left(\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}\right|^{p}\right]\right)^{1/p}\left(\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|Y^{\theta}\right|^{q}\right]\right)^{1/q}=\|\Phi(X)\|_{\mathcal{P}_{p}}\|\Phi(Y)\|_{\mathcal{P}_{q}}$
Minkowski Inequality Parametric Derivation:
  1. 1.
    Translation: $\left(\hat{\mathbb{E}}\left[|X+Y|^{p}\right]\right)^{1/p}=\left(\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}+Y^{\theta}\right|^{p}\right]\right)^{1/p}$.
  2. 2.
    Linear Execution: For each fixed $\theta$, apply the classical Minkowski inequality: $\left(\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}+Y^{\theta}\right|^{p}\right]\right)^{1/p}\leq\left(\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}\right|^{p}\right]\right)^{1/p}+\left(\mathbb{E}_{\mathbb{P}}\left[\left|Y^{\theta}\right|^{p}\right]\right)^{1/p}$.
  3. 3.
    Uniform Synthesis: Taking the supremum over $\theta$ preserves the inequality due to the subadditivity of the supremum operation:
    $\left(\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}+Y^{\theta}\right|^{p}\right]\right)^{1/p}\leq\left(\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}\right|^{p}\right]\right)^{1/p}+\left(\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|Y^{\theta}\right|^{p}\right]\right)^{1/p}$
    This yields the $G$-Minkowski inequality: $\|X+Y\|_{L_{G}^{p}}\leq\|X\|_{L_{G}^{p}}+\|Y\|_{L_{G}^{p}}$.
Burkholder-Davis-Gundy Inequality for $G$-Stochastic Integrals
Classical Statement: For a $G$-Brownian motion $B_{t}$ and an adapted process $H_{t}$, the stochastic integral $\int_{0}^{t}H_{s}dB_{s}$ satisfies for $p>0$:
$c_{p}\hat{\mathbb{E}}\left[\left(\int_{0}^{t}\left|H_{s}\right|^{2}d\langle B\rangle_{s}\right)^{p/2}\right]\leq\hat{\mathbb{E}}\left[\left|\int_{0}^{t}H_{s}dB_{s}\right|^{p}\right]\leq C_{p}\hat{\mathbb{E}}\left[\left(\int_{0}^{t}\left|H_{s}\right|^{2}d\langle B\rangle_{s}\right)^{p/2}\right]$
where $c_{p},C_{p}>0$ are generic constants depending only on $p$.
Parametric Derivation:
  1. 1.
    Translation: By the homomorphism $\Phi$ regarding Itô integrals and quadratic variations:
    $\hat{\mathbb{E}}\left[\left|\int_{0}^{t}H_{s}dB_{s}\right|^{p}\right]=\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\left|\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}\right|^{p}\right]$
    and
    $\hat{\mathbb{E}}\left[\left(\int_{0}^{t}\left|H_{s}\right|^{2}d\langle B\rangle_{s}\right)^{p/2}\right]=\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\left(\int_{0}^{t}\left|H_{s}^{\theta}\right|^{2}\theta_{s}^{2}ds\right)^{p/2}\right]$
  2. 2.
    Linear Execution: For each fixed $\theta$, the process $M_{t}^{\theta}=\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}$ is a classical martingale with quadratic variation $\langle M^{\theta}\rangle_{t}=\int_{0}^{t}\left|H_{s}^{\theta}\theta_{s}\right|^{2}ds=\int_{0}^{t}\left|H_{s}^{\theta}\right|^{2}\theta_{s}^{2}ds$. The classical BDG inequality applies:
    $c_{p}\mathbb{E}_{\mathbb{P}}\left[\langle M^{\theta}\rangle_{t}^{p/2}\right]\leq\mathbb{E}_{\mathbb{P}}\left[|M_{t}^{\theta}|^{p}\right]\leq C_{p}\mathbb{E}_{\mathbb{P}}\left[\langle M^{\theta}\rangle_{t}^{p/2}\right]$
  3. 3.
    Uniform Synthesis: Crucially, the constants $c_{p}$ and $C_{p}$ in the classical BDG inequality depend only on $p$ and are independent of the underlying probability measure or the specific filtration, thus they are independent of $\theta$. Taking the supremum over $\theta\in\mathcal{A}^{\Theta}$ on all sides yields the result directly. For the upper bound:
    $\hat{\mathbb{E}}\left[\left|\int_{0}^{t}H_{s}dB_{s}\right|^{p}\right]$ $=\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|M_{t}^{\theta}\right|^{p}\right]$
    $\leq\sup_{\theta}\left(C_{p}\mathbb{E}_{\mathbb{P}}\left[\left(\int_{0}^{t}\left|H_{s}^{\theta}\right|^{2}\theta_{s}^{2}ds\right)^{p/2}\right]\right)$
    $=C_{p}\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left(\int_{0}^{t}\left|H_{s}^{\theta}\right|^{2}\theta_{s}^{2}ds\right)^{p/2}\right]$
    $=C_{p}\hat{\mathbb{E}}\left[\left(\int_{0}^{t}\left|H_{s}\right|^{2}d\langle B\rangle_{s}\right)^{p/2}\right]$
    The lower bound follows identically.
8.2 Extensions to Fine Analysis and Future Directions
The parametric framework establishes a natural pathway to incorporate Malliavin calculus techniques into the analysis of $G$-expectations. This integration enables a refined sensitivity analysis by working parameter-wise within the Malliavin framework. By examining the Malliavin derivatives and other fine properties for each parameterized classical process $X^{\theta}$, our approach provides a systematic methodology for conducting robust sensitivity analysis under the $G$-expectation framework. This direction opens promising avenues for studying the regularity and sensitivity of solutions to stochastic systems in the presence of volatility uncertainty, which are currently difficult to access using traditional sublinear approaches.
@@ 9 Conclusion
This work has developed a parametric representation framework for $G$-expectations that provides a systematic methodology for stochastic analysis under model uncertainty. We have demonstrated that this approach enables a structured decomposition of complex nonlinear problems, maintaining mathematical rigor while enhancing practical applicability.
The principal advantages of our approach can be summarized as follows:
  • Analytical Efficiency: The clear separation of the computational process into distinct phases allows researchers to leverage the full power of classical stochastic analysis. This structure focuses the specific challenges of nonlinear expectations primarily on the establishment of uniform estimates, thereby streamlining the overall analysis.
  • Enhanced Precision: By performing a unified norm estimation over the entire parameter space at the final stage, our method captures the interdependencies among parameters. This global approach has been shown to yield sharper results compared to methods relying on the composition of piecewise sublinear estimates.
  • Conceptual Clarity: The distinction between linear computational operations and the final robust synthesis offers an intuitive understanding of complex nonlinear phenomena through the lens of parameterized classical processes, making the $G$-framework more accessible.
  • Financial Applications: In mathematical finance, our framework provides a natural foundation for robust pricing and hedging under volatility uncertainty. The parametric representation aligns with stochastic control approaches used in portfolio optimization and derivative pricing, where different volatility scenarios correspond to alternative market models.
Our parametric methodology offers remarkable analytical convenience for studying solution properties. Once well-posedness is established, various characteristics of solutions can be directly investigated through the parameterized family, with complexity concentrated only in the final supremum operation. This not only significantly reduces analytical complexity but also improves estimation precision.
Moreover, the framework exhibits strong extensibility to other types of model uncertainty beyond volatility ambiguity. For instance, in the context of $G$-Lévy processes, our parametric approach can be similarly applied by constructing appropriate parameterized families of classical Lévy processes. This demonstrates the versatility of our methodology in handling diverse forms of Knightian uncertainty within a unified analytical framework.
Building upon the established foundation of $G$-expectation theory, our parametric framework provides a complementary perspective that broadens its practical utility. The results demonstrate that a careful structuring of parameter dependence can yield both theoretical insights and computational benefits.
Looking forward, this parametric approach opens several promising directions for future research. The framework establishes a bridge to computational implementation by reformulating nonlinear problems into tractable families of classical processes. This creates significant potential for developing efficient numerical schemes to compute $G$-expectations and solve related stochastic equations. Furthermore, connections with other advanced stochastic theories and applications in domains such as financial mathematics and engineering systems represent fertile ground for further investigation.
In summary, this work contributes to the development of $G$-expectation theory by providing a structured framework that effectively combines the strengths of classical stochastic analysis with the robustness requirements of model uncertainty. It is our hope that this perspective will facilitate both theoretical advances and practical applications in this evolving field.